OpenStudy (anonymous):

What is the weight (in grams) of a liquid that exactly fills a 163.5 milliliter container if the density of the liquid is 0.492 grams/milliliter ? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (nirmalnema):

density= mass/volume 0.492=mass/163.5 mass=0.492*163.5 mass =80.448gram

OpenStudy (anonymous):

whats it rounded to the nearest hundredth?

OpenStudy (wolf1728):

163.5 ml * .492 gms/ml = 80.44 grams

OpenStudy (anonymous):

thanks!! do you know this one too?

OpenStudy (anonymous):

A large emerald with a mass of 982.7 grams was recently discovered in a mine. If the density of the emerald is 5.42 grams/cm2 , what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (wolf1728):

the density should be 5.42 grams/cm3 shouldn't it?

OpenStudy (wolf1728):

982.7 * 5.42 gms / cm3 = 5,326.23 cm3

OpenStudy (nirmalnema):

density= mass/volume 5.42=982.7/volume volume=982.7/5.42 181.30 cm^3 because it is volume

OpenStudy (anonymous):

is that to the nearest hundredth?

OpenStudy (nirmalnema):

@wolf1728 it should be divided not multiplied

OpenStudy (anonymous):

thank you both!!

OpenStudy (wolf1728):

yes it isYes you are right mirmalnema

OpenStudy (nirmalnema):

yup....

OpenStudy (wolf1728):

Acvtually it should be 181.31 because the exact figure is 181.3099630996